OpenStudy (anonymous):

When A is changing at a rate of -0.120M⋅s−1 , how fast is B changing?

OpenStudy (cuanchi):

depends of the stoichiometry

OpenStudy (cuanchi):

aA+bB-> cC+dD \[Rate= -\frac{ \Delta[A] }{a \Delta t }=-\frac{ \Delta[B] }{b \Delta t }=+\frac{ \Delta[C] }{c \Delta t }=+\frac{ \Delta[D] }{d \Delta t }\]